A concert loudspeaker suspended high off the ground emits 35.0W of sound power. A small microphone with a 0.800cm2 area is 53.0m from the speaker.

How much sound energy impinges on the microphone each second?
Please,show me.
Answer #1

intensity at microphone, I = P/A = 35/(4*pi*r^2) = 35/(4*3.14*53^2) = 9.92*10^-4 W/m^2


power through microphone = A*I = 0.8*10^-4*9.92*10^-4 = 7.936*10^-8 watts


> just change the unit of the second answer to joules and everything else is right

Mary E Ronan Sat, Nov 20, 2021 2:49 PM